# Impact of domain reduction techniques in polynomial optimization: A computational study

### v1

@@ Abstract
Domain reduction techniques are at the core of any global optimization solver for NLP or MINLP problems. In this paper, we delve into several of these techniques and assess the impact they may have in the performance of an RLT-based algorithm for polynomial optimization problems. These techniques include i) the use of (nonlinear) conic relaxations for optimality-based bound tightening, ii) the use of Lagrangian dual information to enhance feasibility-based bound tightening, and iii) different strategies for branching point selection. We also explore how a solver equipped with these domain reduction enhancements can further improve its performance by using machine learning to better choose the best domain reduction approach to use on a given instance.
@@ 1 Introduction
The design and implementation of global optimization algorithms for general MINLP problems is a very active field of research, and the number of available solvers has been steadily increasing over the past years. Recently, companies behind state-of-the-art MILP solvers such as Xpress [xpress] and Gurobi [gurobi] have announced the release of new versions capable of solving general MINLP problems to certified global optimality. Convexifications and domain reduction techniques are probably the two most important elements behind the efficiency of the spatial branching required to handle the nonconvexities as can be seen, for instance, in [Ryoo:1996], [Tawarmalani2004], and [Belotti2009]. The focus of this paper is precisely on the joint assessment of the individual impact of different aspects of domain reduction on the performance of a global optimization algorithm. We hope our analysis can guide the efforts in the development and implementation of this type of enhancements in present and future global optimization solvers for nonconvex problems.
@@
Two essential aspects of domain reduction, present in any spatial branching algorithm, are the selection of the branching variable and the selection of the branching point, since both of them are done at each and every node of the branch-and-bound tree. The choice of the branching variable is relatively well understood, partially because of the thorough research for MILP problems (refer, for instance, to the seminal works [Linderoth1999] and [Achterberg2005]) but also due to some specialized papers for spatial branching such as [Tawarmalani2004] and [Belotti2009]. In the specific context of polynomial optimization, the impact of the criterion for variable selection is thoroughly discussed in [Gonzalez-Rodriguez:2023] and [Ghaddar:2023]. On the other hand, the impact of the selection of the branching point has received much less attention and, although different approaches are mentioned in [Ryoo:1996], [Tawarmalani2004], and [Belotti2009], we are not aware of any numerical analysis assessing how relevant these decisions may be in the performance of the resulting algorithm. In order to bridge this gap, one of the domain reduction aspects we cover in this paper is precisely the selection of the branching point.
@@
Despite the relevance of the selection of both the branching variable and the branching point, probably the single most important aspect in domain reduction in nonlinear programming, at least in terms of the research it has generated, is bound tightening [Ryoo:1996, Tawarmalani2004, Belotti2009, Puranik:2017aa]. Bound tightening techniques are a family of methods designed to reduce the search space by adjusting the bounds of the variables. These methods are often divided into i) optimality-based bound tightening, OBBT, in which bounds are tightened by solving a series of relaxations of minor variations of the original problem and ii) feasibility-based bound tightening, FBBT, in which tighter bounds are deduced directly by exploiting the relations between the problem constraints and the variable bounds. Since OBBT is much more demanding computationally than FBBT, combined schemes are normally used, under which OBBT is only performed at selected nodes (often just at the root node) and FBBT is performed at every node of the branch-and-bound tree. In [Puranik:2017aa], the authors present a thorough computational analysis on the huge impact that domain reduction techniques, and bound tightening in particular, have in three state-of-the-art global optimizers for NLP and MINLP problems. [Gonzalez-Rodriguez:2023] reports a similar impact in the context of an RLT-based algorithm for continuous polynomial optimization problems.
@@
Although integrating a basic bound tightening scheme in a global optimization solver is not too demanding, there is a wide range of sophisticated approaches that have been discussed in the literature and whose implementation may be significantly more complex. [Belotti2013] proposes to enhance FBBT by using convex combinations of constraints instead of only the original ones, while [Ryoo:1996] and [Gleixner:2017] present FBBT enhancements that use Lagrangian dual information from subproblems previously solved by the algorithm. The latter two approaches can be seen as particular cases of the PU-PR domain reduction framework presented in [Tawarmalani2004]. These enhanced methods often entail an additional computational overhead node by node, which should then be overcome by the reduction in the size of the resulting branch-and-bound tree. A second contribution of this paper is devoted to analyze this trade-off, by studying the impact of these sophisticated approaches when integrated into a solver that already has a basic bound tightening scheme in place. More precisely, we study, in the context of the RLT-technique for polynomial optimization [Sherali:1992aa], the impact of some of the Lagrangian-based enhancements in [Ryoo:1996] and [Gleixner:2017].
@@
The standard implementations of optimality-based bound tightening require to solve two subproblems for each problem variable, one to adjust its lower bound and one to adjust its upper bound. Moreover, these problems are typically of a difficulty comparable to that of the nodes solved during the branching process itself. A novel contribution of this paper is to study the use of (nonlinear) conic relaxations, in particular second-order cone programming (SOCP) and semidefinite programming (SDP) ones, to improve the tightening obtained by OBBT. This approach is motivated by the recent analysis in [Gonzalez-Rodriguez:2022] where, in the context of polynomial optimization problems, the authors show the potential of using (nonlinear) conic constraints to tighten the (linear) relaxations solved by an RLT-based algorithm. We adapt their ideas with the goal of improving the performance of the underlying bound tightening techniques.
@@
It is worth noting that domain reduction techniques are not only relevant for theoretical purposes or for the design of general purpose solvers. There have also been important efforts to adapt these techniques to particular classes of problems, directly connected with practical applications. Particularly important are those related to optimization of power networks in general [coffrin2015] and, more specifically, to the alternating current optimal power flow problem or ACOPF, a very active field of research. [Chen:2016] shows the effectiveness of domain reduction techniques for ACOPF problems, [Sundar2018] presents an optimality-based bound tightening to improve the bounds in different parts of the network, using a (convex) quadratic relaxation of the problem, and [Shchetinin2019] proposes three new methods for reducing the domain, focusing on being efficient on large-scale grids.
@@ 2.1 Polynomial optimization, RLT and bound tightening
where is a multiset containing the information about the multiplicity of each variable in the underlying monomial. This linear relaxation is then solved at each node of the branch-and-bound tree. In order to get tighter relaxations and ensure convergence, additional constraints, called bound-factor constraints, are also added (and linearized). These constraints are given, for each pair of multisets and such that , by
In [Dalkiran:2013aa] the authors identify a collection of monomials, which they call -sets, such that convergence to a global optimum is still guaranteed if only the bound-factor constraints associated with these monomials are considered. We adopt this approach since it can dramatically reduce the size of the resulting relaxation (at the price of getting slightly looser relaxations).
@@
OBBT techniques typically solve two optimization problems for each variable, just by substituting the objective function with the variable at hand and then minimizing and maximizing the resulting optimization problem. This procedure gives the best possible bounds for the variable given the constraints and the bounds of the other variables. The downside is that these subproblems may be as hard to solve as the original problem, so the process is often computationally prohibitive. Because of this, OBBT is normally applied to relaxations of the original problem, easier to solve, or letting it run only for a limited amount on time. In the baseline OBBT we consider we do both things: it runs on the RLT linear relaxations and does it for a limited amount of time relative to the total time available.
Even if relaxations are used for OBBT, it can still be computationally demanding, which motivates that OBBT is often used only at the root node (the approach we take) or at a few selected points of the algorithm. FBBT, on the other hand, uses constraint propagation techniques to tighten the bounds and, since no optimization problem is solved in the process, the computational cost is typically so small that FBBT is usually applied at each at every node of the branch-and-bound tree. This is precisely what we do, with an FBBT implementation that builds upon [Belotti:2012aa], where interval arithmetic and a two phase process are used to infer bounds on the variables.
@@ 2.2 Testing environment
For the numerical results presented in this paper, we use the polynomial optimization solver RAPOSa [Gonzalez-Rodriguez:2023], whose core is an RLT-based algorithm. We use three sets of instances to test the performance of the various domain reduction techniques under consideration. The first one is taken from [Dalkiran:2016aa] and consists of 180 instances of randomly generated polynomial optimization problems of different degrees, number of variables, and density. The second test set comes from the well-known benchmark MINLPLib [Bussieck:2003aa], a library of mixed-integer nonlinear programming instances. We have selected from MINLPLib those instances that are (PO) problems with box-constrained and continuous variables, resulting in a total of 166 instances. The third test set comes from another well-known benchmark, QPLIB [Furini:2018aa], a library of quadratic programming instances, for which we made a selection analogous to the one made for MINLPLib, resulting in a total of 63 instances. Then, we removed 6 additional instances that lead to numerical issues in some parts of the analysis, leading to a final set of 403 instances.
@@
To present the results we use different measures to assess the performance of the different approaches. We describe them here to avoid repetition and simplify the notation:
@@
Gap. Geometric mean of the optimality gap obtained by each approach. We disregard instances for which i) at least one approach did not return an optimality gap and (ii) all the approaches solved it within the time limit.
@@
Pace. Geometric mean of pace, as introduced in [Ghaddar:2023]. It represents the number of seconds needed to improve the lower bound of the algorithm by one unit (the pace at which the lower bound improved). We disregard instances solved by every approach in less than 5 seconds. As thoroughly discussed in [Ghaddar:2023], the main motivation behind this performance measure is that it allows to compare the performance on all instances together, whereas Time and Gap fail to do so. Time is not informative when comparing performance between instances not solved by any approach (all of them reach the time limit). These instances can be compared using Gap which, in turn, is not informative in instances solved by all approaches (all of them close the gap), and where Time could be used. Pace, on the other hand, is informative regardless of the number of configurations that might have solved each of the different instances.
@@
We have chosen to report geometric means in our results, with the exception of Bounds and Time, for which the standard mean is used. Geometric means are known to have the advantage of being less sensitive to outliers, which often makes them preferable for performance comparisons. The reason to report the standard mean for Bounds is that it measures how much the bounds are tightened with the different approaches, but it is not really a performance measure as Gap, Time, or Pace, for instance. Thus, since Bounds measures how much certain parameters of the problem are changed before the branch and bound starts, reporting the standard averages allows to get a better picture of the impact of the bound tightening phase on the search space of the algorithm. Similarly, Time is not a measure of the overall performance of a given approach, but just a measure of the time spent on bound tightening at the root node. Reporting standard averages provides a more informative quantification of the impact of a given approach during this phase of the algorithm.111Note that we do not report the number instances for which each approach did not return an optimality gap. The reason for this is that there are not significant differences across approaches and so disregarding them simplifies the tables of results.
@@ 3.1 OBBT enhancements based on conic optimization
SDP: For each -set, we define the vector and we build matrix . We denote the linearization of matrix by , obtained by replacing all monomials in it by the corresponding RLT variables. Then, we add, to the linear relaxation solved at the OBBT phase, the constraint (positive semidefiniteness).
SDP: We do the same as in SDP but changing the vector to .
@@ 3.1.1 Numerical results
SOCP: We follow approach SOCP above and solve the relaxations with Gurobi.
SOCP: We follow approach SOCP above and solve the relaxations with Mosek.
SDP: We follow approach SDP above and solve the relaxations with Mosek.
SDP: We follow approach SDP above and solve the relaxations with Mosek.
@@
SOCP | 285 | 0.133 | 97.67 | 9.98 | 104.38 | 0.341 | 391.92 |
SOCP | 286 | 0.128 | 65.69 | 6.99 | 102.20 | 0.344 | 287.92 |
SDP | 286 | 0.150 | 114.42 | 15.19 | 103.62 | 0.339 | 404.08 |
SDP | 286 | 0.159 | 132.93 | 17.89 | 103.63 | 0.345 | 433.57 |
In Table 1 we present the results associated to the aforementioned approaches, where shaded numbers are used to highlight whenever an approach outperforms the baseline for a certain metric. At first sight, it seems that the overall performance of the new conic approaches for OBBT is worse than the one delivered by Baseline, specially for the SDP ones. As expected, with respect to Baseline, all the conic approaches reduce the number of nodes of the branch and bound tree and improve the tightness of the bounds (Bounds). Yet, the computational overhead required by OBBT at the root node (Time) is not compensated by the modest improvement on Bounds. This is is also captured by the pace of the different approaches: although SOCP and SOCP are very competitive in terms of Gap, the fact that they are more time consuming than Baseline also makes them worse in terms of Pace. SOCP is clearly the best performing conic approach, being fairly competitive with Baseline. Indeed, if we looked at standard means instead of geometric ones (not shown in the table), we would see that SOCP and Baseline are essentially tied for both time ( vs ) and pace ( vs ). The main weakness of the conic approaches, and specially the SDP ones, is that they are more “risky”, in the sense that, since conic solvers are not yet as efficient as linear ones, a conic approach may turn out to be particularly slow in some instances, which might overshadow promising behavior on others.
@@
(403 instances) | 5% Gap (111) | 5% Time (153) | 5% Pace (268) | 5% Nodes (285) | 5% Bounds(403) | 5% Time(403) |
@@
SOCP | 88 | 42 | 150 | 209 | 259 | 68 |
SOCP | 103 | 93 | 200 | 224 | 283 | 141 |
SDP | 82 | 23 | 116 | 214 | 249 | 54 |
SDP | 80 | 10 | 102 | 230 | 358 | 37 |
The results are consistent with those in Table 1, with Baseline and SOCP coming out on top. Importantly, Table 2 also shows that Baseline is not within 5% of the best performing approach in a significant amount of instances: 10% for Gap, 22% for Time, and 15% for Pace. Thus, we could get significant performance improvements if we were able to anticipate the best approach to run on a given instance. This idea is explored in Section 4, where we embed these approaches in a machine learning framework, and train a model with the goal of predicting what the best approach would be on an unseen instance.

### v2

@@ Abstract
Domain reduction techniques are at the core of any global optimization solver for NLP or MINLP problems. In this paper, we delve into several of these techniques and assess the impact they may have in the performance of an RLT-based algorithm for polynomial optimization problems. These techniques include i) the use of (nonlinear) conic relaxations for optimality-based bound tightening, ii) the use of Lagrangian dual information to enhance feasibility-based bound tightening, and iii) different strategies for branching point selection. One of this paper’s main contributions is providing insights into the relative impact of these techniques with respect to each other, which we hope will guide the efforts to develop and implement this type of enhancements in other solvers. We also explore how a solver equipped with these domain reduction enhancements can further improve its performance by using machine learning to better choose the best domain reduction approach to use on a given instance.
@@ 1 Introduction
The design and implementation of global optimization algorithms for general MINLP problems is a very active field of research, and the number of available solvers has been steadily increasing over the past years. Recently, companies behind state-of-the-art MILP solvers, such as Xpress [xpress] and Gurobi [gurobi], have announced the release of new versions capable of solving general MINLP problems to certified global optimality. Convexifications and domain reduction techniques are probably the two most important elements behind the efficiency of the spatial branching required to handle the nonconvexities as can be seen, for instance, in [Ryoo:1996], [Tawarmalani2004], and [Belotti2009]. The literature is full of different enhancements and variations of the core ideas in these papers and, yet, we are aware of no review or consensus on their efficacy beyond the papers in which they were introduced. The main contribution of this paper is precisely the joint assessment of the individual impact of different aspects of domain reduction on the performance of a global optimization algorithm. We hope our analysis can guide the efforts in the development and implementation of this type of enhancements in present and future global optimization solvers for nonconvex problems.
@@
Two essential elements present in any spatial branching algorithm that, in a broad sense, are within the scope of domain reduction, are the selection of the branching variable and the selection of the branching point: both of them allow to control how the feasible domain is reduced at a given node and, importantly, they are applied at each and every node of the branch-and-bound tree. The choice of the branching variable is relatively well understood, partially because of the thorough research for MILP problems (refer, for instance, to the seminal works [Linderoth1999] and [Achterberg2005]) but also due to some specialized papers for spatial branching such as [Tawarmalani2004] and [Belotti2009]. In the specific context of polynomial optimization, the impact of the criterion for variable selection is thoroughly discussed in [Gonzalez-Rodriguez:2023] and [Ghaddar:2023]. On the other hand, the impact of the selection of the branching point has received much less attention and, although different approaches are mentioned in [Ryoo:1996], [Tawarmalani2004], and [Belotti2009], we are not aware of any numerical analysis assessing how relevant these decisions may be for the performance of the resulting algorithm. In order to bridge this gap, one of the domain reduction aspects we cover in this paper is precisely the selection of the branching point.
@@
Despite the relevance of the selection of both the branching variable and the branching point, probably the single most important aspect in domain reduction in nonlinear programming, at least in terms of the research it has generated, is bound tightening [Ryoo:1996, Tawarmalani2004, Belotti2009, Puranik:2017aa]. Bound tightening techniques are a family of methods designed to reduce the search space by adjusting the bounds of the variables. These methods are often divided into i) optimality-based bound tightening, OBBT, in which bounds are tightened by solving a series of relaxations of minor variations of the original problem and ii) feasibility-based bound tightening, FBBT, in which tighter bounds are deduced directly by exploiting the relations between the problem constraints and the variable bounds. In [Puranik:2017aa], the authors present a thorough computational analysis on the huge impact that domain reduction techniques, and bound tightening in particular, have in three state-of-the-art global optimizers for NLP and MINLP problems. [Gonzalez-Rodriguez:2023] reports a similar impact in the context of an RLT-based algorithm for continuous polynomial optimization problems.
@@
Although integrating a basic bound tightening scheme in a global optimization solver is not too demanding, there is a wide range of sophisticated approaches that have been discussed in the literature and whose implementation may be significantly more complex. [Belotti2013] proposes to enhance FBBT by using convex combinations of constraints instead of only the original ones, while [Ryoo:1996] and [Gleixner:2017] present FBBT enhancements that use Lagrangian dual information from subproblems previously solved by the algorithm. The latter two approaches can be seen as particular cases of the duality-based domain reduction framework presented in [Tawarmalani2004]. These enhanced methods often entail an additional computational overhead node by node, which should then be overcome by the reduction in the size of the resulting branch-and-bound tree. A second contribution of this paper is devoted to analyze this trade-off. We do so by studying the impact of these sophisticated approaches when integrated into a solver that already has a basic bound tightening scheme in place. More precisely, we study, in the context of the RLT-technique for polynomial optimization [Sherali:1992aa], the impact of some of the Lagrangian-based enhancements in [Ryoo:1996] and [Gleixner:2017].
@@
A novel contribution of this paper is to study the use of (nonlinear) conic relaxations, in particular second-order cone programming (SOCP) and semidefinite programming (SDP) ones, to improve the tightening obtained by OBBT. This approach is motivated by the recent analysis in [Gonzalez-Rodriguez:2022], where the authors follow the recent trend of using conic relaxations in the context of polynomial optimization [Jeyakumar2017, bonami2019, elloumi2019, burer2020], and show the potential of using (nonlinear) conic constraints to tighten the (linear) relaxations solved by an RLT-based algorithm. We adapt their ideas with the goal of improving the performance of the underlying bound tightening techniques. The above use of SDP relaxations to solve polynomial optimization problems differs greatly from the approach initiated by J. B. Lasserre in [lasserre2001], with the introduction of semidefinite relaxations corresponding to liftings of the polynomial programs into higher dimensions. It has been shown that, as one moves up in the SDP hierarchy, one eventually solves the original polynomial optimization problem. Unfortunately, this approach is currently computationally prohibitive in general, since the sizes of the SDP relaxations quickly become unmanageable (refer to [nie2023] for a recent monograph on the topic).
@@
It is worth noting that domain reduction techniques are not only relevant for theoretical purposes or for the design of general purpose solvers. There have also been important efforts to adapt these techniques to particular classes of problems, directly connected with practical applications. Particularly important are those related to optimization of power networks in general [coffrin2015] and, more specifically, to the alternating current optimal power flow problem or ACOPF, a very active field of research. [Chen:2016] shows the effectiveness of domain reduction techniques for ACOPF problems, [Sundar2018] presents an optimality-based bound tightening scheme to improve the bounds in different parts of the network, using a (convex) quadratic relaxation of the problem, and [Shchetinin2019] proposes three new methods for reducing the domain, focusing on being efficient on large-scale grids.
@@ 2.1 Polynomial optimization, RLT and bound tightening
where is a multiset containing the information about the multiplicity of each variable in the underlying monomial. This linear relaxation is then solved at each node of the branch-and-bound tree. In order to get tighter relaxations and ensure convergence, additional constraints, called bound-factor constraints, are also added (and linearized).111[Sherali:1992aa] proves (see Theorem 1) that, with the inclusion of the bound-factor constraints, the RLT-based algorithm converges to a global optimum of the problem. More precisely, “‘either the branch-and-bound algorithm terminates finitely with the incumbent solution being optimal to (PO), or else an infinite sequence of stages is generated such that along any infinite branch of the branch-and-bound tree, any accumulation point of the -variable sequence of the linear programming iterates generated at the nodes solves (PO)”. These constraints are given, for each pair of multisets and such that the monomial associated with belongs to , by
Given that the size of corresponds with , the number of bound-factor constraints is often prohibitively large. In [Dalkiran:2013aa], the authors identify a collection of monomials, which they call -sets, such that convergence to a global optimum is still guaranteed if only the bound-factor constraints associated with these monomials are considered. -sets are defined as those monomials of degree greater than one present in PO which, moreover, are not included in any other monomial of PO (multiset inclusion). We adopt this approach, since it can dramatically reduce the size of the resulting relaxation (at the price of getting slightly looser relaxations).
@@
OBBT techniques typically solve two optimization problems for each variable, just by substituting the objective function with the variable at hand and then minimizing and maximizing the resulting optimization problem. This procedure gives the best possible bounds for the variable given the constraints and the bounds of the other variables. The downside is that these subproblems may be as hard to solve as the original problem, so the process is often computationally prohibitive. Because of this, OBBT is normally applied to relaxations of the original problem, easier to solve, or letting it run only for a limited amount on time. Our baseline OBBT does both things: it runs on the RLT linear relaxations and does it for a limited amount of time relative to the total time available.
Even if relaxations are used for OBBT, it can still be computationally demanding, which motivates that OBBT is often used only at the root node (the approach we take) or at a few selected points of the algorithm. FBBT, on the other hand, uses constraint propagation techniques to tighten the bounds and, since no optimization problem is solved in the process, the computational cost is typically so small that FBBT is usually applied at each and every node of the branch-and-bound tree. This is precisely what we do, with an FBBT implementation that builds upon [Belotti:2012aa], where interval arithmetic and a two phase process are used to infer bounds on the variables.
@@ 2.2 Testing environment
For the numerical results presented in this paper, we use the polynomial optimization solver RAPOSa [Gonzalez-Rodriguez:2023], whose core is an RLT-based algorithm. RAPOSa is freely available at https://raposa.usc.es. We use three sets of instances to test the performance of the various domain reduction techniques under consideration. The first one is taken from [Dalkiran:2016aa] and consists of 180 instances of randomly generated polynomial optimization problems of different degrees, number of variables, and density (proportion of monomials present in the problem with respect to the total number of monomials of degree at most ). The second test set comes from the well-known benchmark MINLPLib [Bussieck:2003aa], a library of mixed-integer nonlinear programming instances. We have selected from MINLPLib those instances that are (PO) problems with box-constrained and continuous variables, resulting in a total of 166 instances. The third test set comes from another well-known benchmark, QPLIB [Furini:2018aa], a library of quadratic programming instances, for which we made a selection analogous to the one made for MINLPLib, resulting in a total of 63 instances. Then, we removed 6 additional instances that lead to numerical issues in some parts of the analysis, leading to a final set of 403 instances. Instances from the three datasets can be downloaded at https://raposa.usc.es/files/DS-MINLPLib-QPLIB.zip.
@@
We use different measures to assess the performance of the different approaches. We describe them here to avoid repetition and simplify the notation:
@@
Gap. Geometric mean of the optimality gap obtained by each approach. We disregard instances for which i) at least one approach did not return an optimality gap and (ii) all the approaches solved it within the time limit.222The choice of the geometric mean for the gap is motivated by the fact that, for some instances, , which means that the relative gap, computed as , can take values arbitrarily large. Also, there is no need to rely on the shifted mean to account for values very close to zero since, for the numerical analysis, the value of the gap for all solved instances is set to , the stopping tolerance.
@@
Pace. Geometric mean of , as introduced in [Ghaddar:2023]. It represents the number of seconds needed to improve the lower bound of the algorithm by one unit (the pace at which the lower bound improved). We disregard instances solved by every approach in less than 5 seconds. As thoroughly discussed in [Ghaddar:2023], the main motivation behind this performance measure is that it allows to compare the performance on all instances together, whereas Time and Gap fail to do so. Time is not informative when comparing performance between instances not solved by any approach (all of them reach the time limit). These instances can be compared using Gap which, in turn, is not informative in instances solved by all approaches (all of them close the gap), and where Time could be used. Pace, on the other hand, is informative regardless of the number of configurations that might have solved each of the different instances.
@@
We have chosen to report geometric means in our results, with the exception of Bounds and Time, for which the standard mean is used. Geometric means are known to have the advantage of being less sensitive to outliers, which often makes them preferable for performance comparisons. The reason to report the standard mean for Bounds is that it measures how much the bounds are tightened with the different approaches, but it is not really a performance measure as Gap, Time, or Pace, for instance. Thus, since Bounds measures how much certain parameters of the problem are changed before the branching starts, reporting the standard averages allows to get a better picture of the impact of the bound tightening phase on the search space of the algorithm. Similarly, Time is not a measure of the overall performance of a given approach, but just a measure of the time spent on bound tightening at the root node. Reporting standard averages provides a more informative quantification of the impact of a given approach during this phase of the algorithm.333Note that we do not report the number of instances for which each approach did not return an optimality gap. The reason for this is that there are not significant differences across approaches and so disregarding them simplifies the tables of results.
@@ 3.1 OBBT enhancements based on conic optimization
SDP1: For each -set, we define the vector and we build matrix . We denote the linearization of matrix by , obtained by replacing all monomials in it by the corresponding RLT variables. Then, we add, to the linear relaxation solved at the OBBT phase, the constraint (positive semidefiniteness).
SDP2: We do the same as in SDP1 but changing the vector to .
@@ 3.1.1 Numerical results
Baseline: We use the standard OBBT and FBBT used in [Gonzalez-Rodriguez:2023]. At the root node, we perform OBBT before solving it. OBBT consists of minimizing and maximizing each variable over the feasible set of the root node (i.e., its linear relaxation). Moreover, before solving each node, we perform FBBT using interval arithmetic.
SOCPG: We follow approach SOCP above and solve the relaxations with Gurobi.
SOCPM: We follow approach SOCP above and solve the relaxations with Mosek.
SDP1: We follow approach SDP1 above and solve the relaxations with Mosek.
SDP2: We follow approach SDP2 above and solve the relaxations with Mosek.
@@
| SOCPG | 285 | 0.133 | 97.67 | 9.98 | 104.38 | 0.341 | 391.92 |
| SOCPM | 286 | 0.128 | 65.69 | 6.99 | 102.20 | 0.344 | 287.92 |
| SDP1 | 286 | 0.150 | 114.42 | 15.19 | 103.62 | 0.339 | 404.08 |
| SDP2 | 286 | 0.159 | 132.93 | 17.89 | 103.63 | 0.345 | 433.57 |
In Table 1 we present the results associated to the aforementioned approaches, where shaded numbers are used to highlight whenever an approach outperforms the baseline for a certain metric. At first sight, it seems that the overall performance of the new conic approaches for OBBT is worse than the one delivered by Baseline, specially for the SDP ones. As expected, with respect to Baseline, all the conic approaches reduce the number of nodes of the branch-and-bound tree and improve the tightness of the bounds (Bounds). Yet, the computational overhead required by OBBT at the root node (Time) is not compensated by the modest improvement on Bounds. This is is also captured by the pace of the different approaches: although SOCPG and SOCPM are very competitive in terms of Gap, the fact that they are more demanding computationally than Baseline also makes them worse in terms of Pace. SOCPM is clearly the best performing conic approach, being fairly competitive with Baseline. Indeed, if we looked at standard means instead of geometric ones (not shown in the table), we would see that SOCPM and Baseline are essentially tied for both time ( vs ) and pace ( vs ). The main weakness of the conic approaches, and specially the SDP ones, is that they are more “risky”, in the sense that, since conic solvers are not yet as efficient as linear ones, a conic approach may turn out to be particularly slow in some instances, which might overshadow promising behavior on others.
@@
| (403 instances) | Gap5% (111) | Time5% (153) | Pace5% (268) | Nodes5% (285) | Bounds (403) | Time (403) |
@@
| SOCPG | 88 | 42 | 150 | 209 | 259 | 68 |
| SOCPM | 103 | 93 | 200 | 224 | 283 | 141 |
| SDP1 | 82 | 23 | 116 | 214 | 249 | 54 |
| SDP2 | 80 | 10 | 102 | 230 | 358 | 37 |
The results are consistent with those in Table 1, with Baseline and SOCPM coming out on top. Importantly, Table 2 also shows that Baseline is not within 5% of the best performing approach in a significant amount of instances: 10% for Gap, 22% for Time, and 15% for Pace. Thus, we could get significant performance improvements if we were able to anticipate the best approach to run on a given instance. This idea is explored in Section 4, where we embed these approaches in a machine learning framework, and train a model with the goal of predicting what the best approach would be on an unseen instance.
Finally, it is worth noting that, in the analysis carried out in [Gonzalez-Rodriguez:2022], the authors managed to successfully identify families of instances where one specific approach was particularly effective. Unfortunately, we have not been able to find such patterns behind the results in Table 1 and Table 2, or any other set of results presented in this work.